# Short vs. Long-term Coordination of Drones: When Distributed Optimization Meets Deep Reinforcement Learning

### v3

@@ II Related Work
Notation Explanation
$u,U,\mathcal{U}$ Index of a drone; total number of drones; set of drones
$m,M,\mathcal{M}$ Index of a charging station; total number of charging stations; set of charging stations
$n,N,\mathcal{N}$ Index of a grid cell; total number of grid cells; set of grid cells
$s,S,\mathcal{S}$ Index of a timeslot; total number of timeslots in a period; set of timeslots
$t,T,\mathcal{T}$ Index of a period; total number of periods; set of periods
$a^{u}$ The action or flying direction taken by $u$
$P^{u},p^{u}_{n,s}$ The plan of $u$; value of $P^{u}$ at cell $n$ and timeslot $s$
$P^{-u},p^{-u}_{n,s}$ The observed plans of drones by $u$ excluding $P^{u}$; value of $P^{-u}$ at cell $n$ and timeslot $s$
$p_{n,s}$ The aggregated plans of all drones at cell $n$ and timeslot $s$
$V_{n,s}$ Required sensing value at cell $n$ and timeslot $s$
$v_{n,s}$ Collected sensing value at cell $n$ and timeslot $s$
$\mathcal{R},r_{n,s}$ Target; target value at cell $n$ and timeslot $s$
$l,L$ Index of a plan in $G^{a^{u}}$; total number of plans in $G^{a^{u}}$
$C^{\mathsf{f}}(u),t^{\mathsf{f}}$ Flying power consumption of $u$; flying time
$C^{\mathsf{h}}(u),t^{\mathsf{h}}_{u}$ Hovering power consumption of $u$; hovering time
$c^{u},e^{u}$ Current location of $u$; current energy consumption of $u$
$K(a^{u})$ Indexes of cells within the searching range along flying direction $a^{u}$
$J(a^{u})$ Indexes of visited cells within $K(a^{u})$
$\beta$ Behavior of an agent in planning optimization
$o^{u}_{t},r^{u}_{t}$ Observation of $u$ at period $t$; reward of $u$ at the period $t$
$k,H$ Index of a sampled transition; number of sampled transitions
$\gamma,\epsilon$ Discount factor; clip interval hyperparameter
$Q(\cdot),\theta^{Q}$ Critic network; parameter of critic network
$\pi(\cdot),\theta^{\pi}$ Actor network; parameter of actor network
$\omega_{n,s},\overline{v}$ Prediction coefficient to calculate $\hat{V}_{n,s}$; threshold to remove the regions with low sensing requirements
$I,\mathcal{X},A$ Total number of iterations in EPOS; size of state space; number of actions in DRL
$\mathcal{E},W$ Number of episodes; number of neutrals per layer in DNN

### v4

@@ II Related Work
Notation
Explanation
$u,U,\mathcal{U}$
Index of a drone; total number of drones; set of drones
$m,M,\mathcal{M}$
Index of a charging station; total number of charging stations; set of charging stations
$n,N,\mathcal{N}$
Index of a grid cell; total number of grid cells; set of grid cells
$s,S,\mathcal{S}$
Index of a timeslot; total number of timeslots in a period; set of timeslots
$t,T,\mathcal{T}$
Index of a period; total number of periods; set of periods
$a^{u}$
The action or flying direction taken by $u$
$P^{u},p^{u}_{n,s}$
The plan of $u$; value of $P^{u}$ at cell $n$ and timeslot $s$
$P^{-u},p^{-u}_{n,s}$
The observed plans of drones by $u$ excluding $P^{u}$; value of $P^{-u}$ at cell $n$ and timeslot $s$
$p_{n,s}$
The aggregated plans of all drones at cell $n$ and timeslot $s$
$V_{n,s}$
Required sensing value at cell $n$ and timeslot $s$
$v_{n,s}$
Collected sensing value at cell $n$ and timeslot $s$
$\mathcal{R},r_{n,s}$
Target; target value at cell $n$ and timeslot $s$
$l,L$
Index of a plan in $G^{a^{u}}$; total number of plans in $G^{a^{u}}$
$C^{\mathsf{f}}(u),t^{\mathsf{f}}$
Flying power consumption of $u$; flying time
$C^{\mathsf{h}}(u),t^{\mathsf{h}}_{u}$
Hovering power consumption of $u$; hovering time
$c^{u},e^{u}$
Current location of $u$; current energy consumption of $u$
$K(a^{u})$
Indexes of cells within the searching range along flying direction $a^{u}$
$J(a^{u})$
Indexes of visited cells within $K(a^{u})$
$\beta$
Behavior of an agent in planning optimization
$o^{u}_{t},r^{u}_{t}$
Observation of $u$ at period $t$; reward of $u$ at the period $t$
$k,H$
Index of a sampled transition; number of sampled transitions
$\gamma,\epsilon$
Discount factor; clip interval hyperparameter
$Q(\cdot),\theta^{Q}$
Critic network; parameter of critic network
$\pi(\cdot),\theta^{\pi}$
Actor network; parameter of actor network
$\omega_{n,s},\overline{v}$
Prediction coefficient to calculate $\hat{V}_{n,s}$; threshold to remove the regions with low sensing requirements
$I,\mathcal{X},A$
Total number of iterations in EPOS; size of state space; number of actions in DRL
$\mathcal{E},W$
Number of episodes; number of neutrals per layer in DNN